Solve the cash flow equivalence below for the unknown value of Q assuming an 7% annual interest rate. 1. 800 4i Q 70 o0 i23 2 2, You borrowed $6,000.00 for 5 years at 7% annual interest rate. The banker said that to repay the total loan amount you have to pay $1,463 at the end of each year. a) Draw a time line depicting this cash low b) Build a table to determine how much of the annual payment is interest, and 2 how much principal is there in each annual payment. 4 You borrowed $4,000.00 at 1% per month and agreed to repay in equal monthly payments over the next 3 years. What is your monthly payment? 3. 4. You borrowed $10,000.00 which is to be repaid in equal quarterly payments of $1,336.00 over the next 2 years. a) Determine the interest rate per interest period based on quarterly compounding b) What are the nominal and effective interest rates? 2 5. You bought a used car for $4,000.00 at a nominal interest rate of 6%. You agreed to pay for the car in 12 equal monthly payments, beginning with the first payment at the time of the purchase of the car. a) What is the monthly payment? Immediately after making the fifth payment, you made an arrangement with the company to pay back the rest of the loan with one single payment at the time when the sixth payment was due. What is your sixth payment? b)
Answer #1

(1)

Present value of known cash flow = 400 x P/F(7%, 1) + 800 x P/F(7%, 3) = 400 x 0.9346** + 800 x 0.8163**

= 373.84 + 653.04

= 1026.88

For equivalence,

4Q + 7Q x P/F(7%, 3) = 1026.88

4Q + 7Q x 0.8163** = 1026.88

4Q + 5.7141Q = 1026.88

Q x (4 + 5.7141) = 1026.88

Q x 9.7141 = 1026.88

Q = 105.71

**From P/F Factor table

NOTE: As per Answering Policy, 1st question is answered.